# arXiv:2603.02009v1[math.AP] 02 Mar 2026

### v1

@@ Abstract.
This paper investigates the critical quintic wave equation in a 3D bounded domain subject to locally distributed Kelvin-Voigt damping. The study tackles two major mathematical challenges: the severe loss of derivatives induced by the localized thermo-viscous dissipation and the aggressive nature of the critical nonlinear term. First, we establish a robust well-posedness theory for arbitrarily large initial data. By shifting the analysis to the frequency space via a Littlewood-Paley decomposition , we employ Bernstein’s inequalities to lift the damping term into an $L^{2}$ framework, allowing Strichartz estimates to be applied flawlessly. In the second part, we prove the uniform exponential stabilization of the energy. To overcome the reduction of the residual to the $H^{-2}$ level caused by the Kelvin-Voigt mechanism, we utilize the microlocal defect measure framework. The core of our stabilization proof relies on combining the critical Strichartz regularity $L^{4}_{t}L^{12}_{x}$ with a sharp Unique Continuation Property (UCP) to close the observability argument. Furthermore, we demonstrate that this microlocal mechanism is perfectly compatible with non-invasive damping geometries of arbitrarily small Lebesgue measure, successfully circumventing the geometric obstruction of trapped rays.
@@ 1.2. The drawback of Galerkin’s method
We would like to take this opportunity to mention that, in the manuscript [9] by the authors of present article, we assumed, without any explicit mention, that Galerkin’s method would work well for small datasets and that, for large datasets, we would need to prove or assume that we have a Shatah and Struwe solution, as done in [11], [20], although we did cite those references. Since our primary focus was on decay, we ended up not mentioning this fact. Therefore, we would like to apologize to the readers for this oversight.
@@ 1.3. Literature Overview
In the presence of Kelvin–Voigt damping, however, the analytical structure changes significantly. The damping operator introduces higher-order spatial effects which interact nontrivially with geometric multipliers, frequently requiring additional regularity assumptions. This phenomenon has been examined in various stabilization settings, including localized and degenerate dissipation mechanisms (see [1]).
@@ 1.4. Contributions of the present article
To overcome this bottleneck and remove the small data restriction, the second part of our work shifts the paradigm from the physical domain to the frequency space, deploying the modern arsenal of Harmonic Analysis. The application of dispersive techniques to boundary value problems possesses a rich and profound history. The seminal works of Smith and Sogge [27] demonstrated how wave dispersion interacts with domains and rough boundaries. Subsequently, the celebrated contributions of the trio Burq, Lebeau, and Planchon (BLP) solidified the use of spectral projectors and Littlewood-Paley theory tailored to Dirichlet boundary conditions, allowing Strichartz estimates to flourish outside of $\mathbb{R}^{n}$. Furthermore, treating the singular damping terms via the Duhamel formulation crucially relies on the foundational machinery of Christ and Kiselev, which elegantly translates homogeneous dispersive bounds into retarded estimates, allowing low-frequency perturbations to be seamlessly absorbed.
@@
  • The High-Frequency Regime: Conversely, at high frequencies, Bernstein’s inequality cannot be applied without yielding unfavorable positive powers of the frequency threshold $N$. Instead, our strategy exploits the finiteness of the initial energy. Since the high-frequency tail of the initial data converges to zero as $N\to\infty$, its contribution to the critical Strichartz norm can be made arbitrarily small. Thus, the high-frequency regime is controlled not by lifting the derivative, but by the asymptotic smallness of the energy tail.
@@ 4.1. Uniqueness of Strong Solutions
where $C=5M\max(1,C_{P})$. Since $w(0)=w_{t}(0)=0$, we have $E_{w}(0)=0$. By Grönwall’s Lemma, $E_{w}(t)\equiv 0$ for all $t\in[0,T]$, which concludes the proof that $u\equiv v$. ∎
@@ 5. Littlewood-Paley Strategy
In the present section the Littlewood–Paley decomposition reveals an intrinsic stabilizing mechanism: the Kelvin–Voigt operator behaves as a spectral smoothing device. Low frequencies absorb derivative losses via Bernstein inequalities, while high frequencies remain dispersive
@@ 5.1. The Philosophy: Frequencies Instead of Space
The central challenge is the Kelvin-Voigt term: $\text{div}(a(x)\nabla u_{t})$. By having two spatial derivatives ($\nabla\cdot\nabla$), this term belongs to the dual space $H^{-1}$. The classical Strichartz estimate for the wave equation does not accept terms in $H^{-1}$ on the right-hand side.
@@
The spectral solution: Instead of slicing the physical domain, we cut the frequency of the wave. The low-frequency spectral projection operator $P_{\leq N}$ possesses the property of Bernstein’s Inequality. It limits the cost of the derivatives, lifting the term from $H^{-1}$ back to $L^{2}$, a space where the Strichartz Theorem works perfectly.
@@ 5.2. Step 1: The Regularized Problem and the Energy
$u_{tt}-\Delta u-\text{div}(a(x)\nabla u_{t})+f_{k}(u)=0$ $\text{in }\Omega\times(0,T),$ (19)
@@
5.3. Step 2: The Spectral Filter ($P_{\leq N}$)
We use the Spectral Theorem for the $\Delta$ operator with Dirichlet boundary conditions, see the Appendix of the present article. Let $\{\lambda_{j}\}$ be the eigenvalues and $\{\phi_{j}\}$ be the orthonormal eigenfunctions. For a threshold $N>0$, we define the Low-Frequency Projector:
$P_{\leq N}u=\sum_{\sqrt{\lambda_{j}}\leq N}(u,\phi_{j})_{L^{2}}\phi_{j}$
This operator commutes with $\Delta$ and $\partial_{t}$, and satisfies Bernstein’s Inequality:
$\|\nabla(P_{\leq N}g)\|_{L^{2}(\Omega)}\leq N\|g\|_{L^{2}(\Omega)}$ (23)
@@ 5.4. Step 3: Low Frequencies and Strichartz
We define the macroscopic part of the wave $w_{N}=P_{\leq N}u$. Applying $P_{\leq N}$ to the PDE (19):
$(w_{N})_{tt}-\Delta w_{N}=-P_{\leq N}(f_{k}(u))+P_{\leq N}\big(\text{div}(a(x)\nabla u_{t})\big)$ (24)
To tame the Kelvin-Voigt (KV) intrusive term, we transfer the derivative using the adjoint property and Bernstein’s inequality:
$\|P_{\leq N}\text{div}(a(x)\nabla u_{t})\|_{L^{2}}$ $=\sup_{\|\varphi\|_{L^{2}}=1}\langle a(x)\nabla u_{t},\nabla(P_{\leq N}\varphi)\rangle$
$\leq\sup_{\|\varphi\|_{L^{2}}=1}\|a(x)\nabla u_{t}\|_{L^{2}}\cdot(N\|\varphi\|_{L^{2}})$
$=N\|a(x)\nabla u_{t}\|_{L^{2}}$
To justify the interaction between the Kelvin–Voigt operator and the spectral projector, we exploit the selfadjointness of $P_{\leq N}$. For any $\phi\in L^{2}(\Omega)$, we write
@@
Integrating by parts in space yields
@@
Using the Bernstein inequality for spectral projectors,
$\|\nabla(P_{\leq N}\phi)\|_{L^{2}(\Omega)}\leq CN\|\phi\|_{L^{2}(\Omega)},$ (27)
we obtain
$\|P_{\leq N}\operatorname{div}(a\nabla u_{t})\|_{L^{2}(\Omega)}\leq CN\|a\nabla u_{t}\|_{L^{2}(\Omega)}.$ (28)
Integrating in $L^{1}(0,T)$ and using the Global Energy (22):
$\|F_{KV}\|_{L^{1}L^{2}}\leq C_{a}N\sqrt{T}E_{0}^{1/2}$
With the right-hand side in $L^{1}L^{2}$, we apply Strichartz to $w_{N}$:
$\|w_{N}\|_{L^{5}L^{10}}\leq C_{S}E_{0}^{1/2}+C_{S}C_{a}N\sqrt{T}E_{0}^{1/2}+C_{1}\|u\|_{L^{5}L^{10}}^{5}$ (29)
@@
where we used the standard Bernstein inequality for spectral projectors associated with the Dirichlet Laplacian (see Blair–Smith–Sogge).
5.5. Step 4: High Frequencies (The Microscopic Fluctuation)
Let $v_{N}=P_{>N}u=u-w_{N}$. To tame the quintic term, we need $\|v_{N}\|_{L^{5}L^{10}}$ to be small. Since the Sobolev embedding $H^{1}\hookrightarrow L^{10}$ fails in 3D, we exploit dispersion via initial data. Because the initial data $(u_{0},u_{1})$ does not depend on the truncation $k$, the projection onto high frequencies goes to zero as $N\to\infty$. Applying the classical Strichartz estimate to the high frequency, we guarantee that for short times:
$\|v_{N}\|_{L^{5}L^{10}}\leq\epsilon(N)$ (30)
where $\epsilon(N)\to 0$ as $N\to\infty$, independently of $k$.

### v2

# The Quintic Wave Equation with Kelvin-Voigt Damping: Strichartz Estimates, Well-posedness and Global Stabilization Revisited

@@ Abstract.
This paper investigates the critical quintic wave equation in a 3D bounded domain subject to locally distributed Kelvin-Voigt damping. The study tackles two major mathematical challenges: the severe loss of derivatives induced by the localized thermo-viscous dissipation and the aggressive nature of the critical nonlinear term. First, we establish a robust well-posedness theory for arbitrarily large initial data. By shifting the analysis to the frequency space via a Littlewood-Paley decomposition , we employ Bernstein’s inequalities to lift the damping term into an $L^{2}$ framework, allowing Strichartz estimates to be applied effectively. In the second part, we prove the uniform exponential stabilization of the energy. To overcome the reduction of the residual to the $H^{-2}$ level caused by the Kelvin-Voigt mechanism, we utilize the microlocal defect measure framework. The core of our stabilization proof relies on combining the critical Strichartz regularity $L^{4}_{t}L^{12}_{x}$ with a sharp Unique Continuation Property (UCP) to close the observability argument. Furthermore, we demonstrate that this microlocal mechanism is perfectly compatible with non-invasive damping geometries of arbitrarily small Lebesgue measure, successfully circumventing the geometric obstruction of trapped rays.
@@ 1.2. The drawback of Galerkin’s method
We would like to emphasize that in the article [9] by the authors of the present article, our main objective was to show that the Strichartz estimates of Blair, Smith, and Sogge [3], together with the unique continuation principle due to Duyckaerts, Zhang, and Zuazua [14], constitute the precise tools for obtaining well-posedeness as well as exponential decay of energy of the quintic wave equation subjected a localized frictional damping $a(x)u_{t}$. For this reason, we assumed, without any explicit mention, that Galerkin’s method would work well for small datasets and that, for large datasets taken in bounded sets of phase space, $H_{0}^{1}\times L^{2}$, we would need to prove (or assume) that we have a Shatah and Struwe solution, as done in [11], [20], although we did cite those references. Since our primary focus was on exponential decay, we did not mention this fact. Therefore, we would like to apologize to the readers for this oversight.
@@ 1.3. Literature Overview
In the presence of Kelvin–Voigt damping, however, the analytical structure changes significantly. The damping operator introduces higher-order spatial effects which interact nontrivially with geometric multipliers, frequently requiring additional regularity assumptions. This phenomenon has been examined in various stabilization settings, including localized and degenerate dissipation mechanisms (see [1]), [24].
@@ 1.4. Contributions of the present article
To overcome this bottleneck and remove the small data restriction, the second part of our work shifts the paradigm from the physical domain to the frequency space, deploying the modern arsenal of Harmonic Analysis. The application of dispersive techniques to boundary value problems possesses a rich and profound history. The seminal works of Smith and Sogge [28] demonstrated how wave dispersion interacts with domains and rough boundaries. Subsequently, the celebrated contributions of the trio Burq, Lebeau, and Planchon [4] (BLP) solidified the use of spectral projectors and Littlewood-Paley theory tailored to Dirichlet boundary conditions, allowing Strichartz estimates to flourish outside of $\mathbb{R}^{n}$. Furthermore, treating the singular damping terms via the Duhamel formulation crucially relies on the foundational machinery of Christ and Kiselev [10], which elegantly translates homogeneous dispersive bounds into retarded estimates, allowing low-frequency perturbations to be seamlessly absorbed.
@@
  • The High-Frequency Regime: Conversely, at high frequencies, Bernstein’s inequality cannot be applied without yielding unfavorable positive powers of the frequency threshold $N$. To circumvent this severe derivative loss, we fundamentally shift our operator strategy by keeping the localized damping within the principal part of the wave operator. Applying the high-frequency projector $P_{>N}$ generates a non-trivial commutator $P_{>N},a(x)]$ with the variable damping coefficient. By leveraging the theory of smooth spectral multipliers and Schur’s test, we rigorously demonstrate that this commutator acts as a smoothing operator of order 1-1. This remarkable cancellation completely absorbs the spatial derivative of the Kelvin-Voigt term without any asymptotic loss. With the damping safely incorporated on the left-hand side, the remaining high-frequency evolution is ultimately controlled by the asymptotic smallness of the initial energy tail (as $N\to\infty$) and a short-time perturbation argument, allowing us to seamlessly close the critical Strichartz estimates.
@@ 4.1. Uniqueness of Strong Solutions
where $C=5M\max(1,C_{P})$. Since $w(0)=w_{t}(0)=0$, we have $E_{w}(0)=0$. By Gronwall’s Lemma, $E_{w}(t)\equiv 0$ for all $t\in[0,T]$, which concludes the proof that $u\equiv v$. ∎
@@ 5. Littlewood-Paley Strategy
In the present section the Littlewood–Paley decomposition reveals an intrinsic stabilizing mechanism: the Kelvin–Voigt operator behaves as a spectral smoothing device. Low frequencies absorb derivative losses via Bernstein inequalities, while high frequencies remain dispersive.
@@ 5.1. The Philosophy: Frequencies Instead of Space
The central challenge is the Kelvin-Voigt term: $\operatorname{div}(a(x)\nabla u_{t})$. By having two spatial derivatives ($\nabla\cdot\nabla$), this term belongs to the dual space $H^{-1}$. The classical non-homogeneous Strichartz estimate for the wave equation does not accept terms in $H^{-1}$ on the right-hand side.
@@
The spectral solution: Instead of slicing the physical domain, we partition the frequency of the wave. To preserve the boundedness of the operators in $L^{p}$ spaces and ensure the convolution kernels decay rapidly, we employ a smooth spectral projection. Let $\chi\in C_{c}^{\infty}(\mathbb{R})$ be a smooth cutoff function such that $\chi(s)=1$ for $|s|\leq 1$ and $\chi(s)=0$ for $|s|\geq 2$. We define the low-frequency projector $P_{\leq N}=\chi(\sqrt{-\Delta}/N)$ and its high-frequency counterpart $P_{>N}=I-P_{\leq N}$. This smooth localization possesses the powerful property of satisfying Bernstein’s Inequalities. It effectively limits the cost of spatial derivatives, lifting the damping term from $H^{-1}$ back to $L^{2}$, which is precisely the functional framework where the Strichartz Theorem operates perfectly.
@@ 5.2. Step 1: The Regularized Problem and the Energy
$u_{tt}-\Delta u-\operatorname{div}(a(x)\nabla u_{t})+f_{k}(u)=0$ $\text{in }\Omega\times(0,T),$ (19)
@@
5.3. Step 2: The Smooth Spectral Filter ($P_{\leq N}$)
We use the Spectral Theorem for the $\Delta$ operator with Dirichlet boundary conditions (see the Appendix of the present article). Let $\{\lambda_{j}\}_{j=1}^{\infty}$ be the eigenvalues and $\{\phi_{j}\}_{j=1}^{\infty}$ be the corresponding orthonormal eigenfunctions.
To ensure the resulting operators have rapidly decaying kernels and bounded commutators, we must employ a smooth spectral cutoff rather than a sharp truncation. For a frequency threshold $N>0$, we define the smooth Low-Frequency Projector via the functional calculus:
$P_{\leq N}u=\sum_{j=1}^{\infty}\chi\left(\frac{\sqrt{\lambda_{j}}}{N}\right)(u,\phi_{j})_{L^{2}}\phi_{j}.$ (23)
By construction, this operator commutes with both $\Delta$ and $\partial_{t}$. Crucially, because $\chi$ is smoothly supported on [2,2][-2,2], the projector rigorously satisfies Bernstein’s Inequality:
$\|\nabla(P_{\leq N}g)\|_{L^{2}(\Omega)}\leq CN\|g\|_{L^{2}(\Omega)},$ (24)
where the constant $C$ depends only on the choice of the profile $\chi$.
@@ 5.4. Step 3: Low Frequencies and Strichartz
We define the macroscopic part of the wave as $w_{N}=P_{\leq N}u$. Applying the spectral projector $P_{\leq N}$ to the truncated PDE (19), we obtain:
$(w_{N})_{tt}-\Delta w_{N}=-P_{\leq N}(f_{k}(u))+P_{\leq N}\big(\operatorname{div}(a(x)\nabla u_{t})\big).$ (25)
To apply the Strichartz estimates to the low-frequency component $w_{N}$, we treat the entire right-hand side of (25) as a source term $F_{total}=F_{NL}+F_{KV}$, where $F_{NL}:=-P_{\leq N}(f_{k}(u))$ and $F_{KV}:=P_{\leq N}\big(\operatorname{div}(a(x)\nabla u_{t})\big)$. Our goal is to estimate both terms in the dual Strichartz space $L^{1}(0,T;L^{2}(\Omega))$.
To justify the interaction between the intrusive Kelvin-Voigt operator and the spectral projector, we exploit the self-adjointness of $P_{\leq N}$. For any test function $\phi\in L^{2}(\Omega)$, we write:
@@
Integrating by parts in space yields:
@@
At this point, we invoke the standard Bernstein inequality for spectral projectors associated with the Dirichlet Laplacian (see, for instance, Blair–Smith–Sogge [3]):
$\|\nabla(P_{\leq N}\phi)\|_{L^{2}(\Omega)}\leq CN\|\phi\|_{L^{2}(\Omega)}.$ (28)
Applying the Cauchy-Schwarz inequality, we deduce that the projector acts as a low-pass filter, recovering the spatial derivative at the cost of the frequency threshold $N$:
$\|F_{KV}(t)\|_{L^{2}(\Omega)}\leq CN\|a\nabla u_{t}(t)\|_{L^{2}(\Omega)}\leq C_{a}NE(t)^{1/2}.$ (29)
Integrating over the time interval $[0,T]$ and using the uniform bound from the first-order global energy $E(t)\leq E_{0}$, we obtain the targeted bound:
$\|F_{KV}\|_{L^{1}(0,T;L^{2}(\Omega))}\leq C_{a}N\sqrt{T}E_{0}^{1/2}.$ (30)
Remark 5.1.
Due to the regularity of the truncated problem we constructed in Section 5, we are not applying the projector $P_{\leq N}$ to the original weak solution. We are applying it to the strong solution $u_{k}$ of the truncated problem! From Section 3, $u_{k}\in H^{2}\cap H_{0}^{1}$ and $\partial_{t}u_{k}\in H_{0}^{1}$. Therefore, $\nabla u_{k,t}\in L^{2}$ and, strictly speaking, $\operatorname{div}(a\nabla u_{k,t})$ would be in $H^{-1}$. However, the low-frequency Littlewood-Paley projector ($P_{\leq N}$) has the wonderful property of being a regularizer. It cuts off the high frequencies, which means that even if you apply $P_{\leq N}$ to a distribution in $H^{-1}$, the result falls across all $H^{s}$ spaces, including $L^{2}$ and $C^{\infty}$.
For the nonlinear term, we use the $L^{p}$-boundedness of the spectral projector. Since the truncated nonlinearity obeys $|f_{k}(s)|\leq|s|^{5}$ uniformly in $k$, we have:
$\|F_{NL}(t)\|_{L^{2}(\Omega)}\leq C\|f_{k}(u(t))\|_{L^{2}(\Omega)}\leq C\||u(t)|^{5}\|_{L^{2}(\Omega)}=C\|u(t)\|_{L^{10}(\Omega)}^{5}.$ (31)
Integrating in time from 0 to $T$ yields exactly the critical Strichartz norm associated with the quintic wave equation:
$\|F_{NL}\|_{L^{1}(0,T;L^{2}(\Omega))}\leq C\|u\|_{L^{5}(0,T;L^{10}(\Omega))}^{5}.$ (32)
Combining (30) and (32), the total source term is bounded in $L^{1}L^{2}$. Applying the standard Strichartz estimate to equation (25), we conclude that for the critical space $X_{T}:=L^{5}(0,T;L^{10}(\Omega))$, the low-frequency component satisfies:
$\|w_{N}\|_{X_{T}}\leq C_{0}E_{0}^{1/2}+C_{1}\|u\|_{X_{T}}^{5}+C_{S}C_{a}N\sqrt{T}E_{0}^{1/2},$ (33)
where $C_{0}$ depends on the free wave propagation of the initial data, $C_{1}$ is a universal constant, and $C_{S}$ is the Strichartz constant.
5.5. Step 4: High Frequencies and the Commutator Trick
Let $v_{N}=P_{>N}u=u-w_{N}$ denote the microscopic fluctuation of the wave. To successfully tame the critical quintic term in our subsequent bootstrap argument, it is mandatory to ensure that the Strichartz norm $\|v_{N}\|_{L^{5}(0,T;L^{10}(\Omega))}$ is strictly small. In dimension $n=3$, the standard Sobolev embedding $H^{1}(\Omega)\hookrightarrow L^{10}(\Omega)$ fails, preventing a direct uniform bound solely from the energy space. To overcome this, we exploit the dispersive decay of the initial data.
Unlike the macroscopic regime (low frequencies), we cannot treat the Kelvin-Voigt operator as a bounded perturbation on the right-hand side of the free wave equation for $v_{N}$. Applying Bernstein’s inequality here would yield a positive power of the frequency threshold $N$, which diverges as $N\to\infty$. Conversely, keeping the term without Bernstein would require bounding $\Delta\partial_{t}u_{k}$ in $L^{2}$, which heavily depends on the truncation parameter $k$ and would destroy the uniformity of the local existence time.
To completely circumvent this loss of derivatives, we keep the localized damping operator on the left-hand side. We define the damped wave operator $L$:
$L:=\partial_{tt}-\Delta-\operatorname{div}(a(x)\nabla\partial_{t}).$ (34)
When we apply the high-frequency spectral projector $P_{>N}$ to the original PDE ($Lu=-f_{k}(u)$), the projector commutes perfectly with the constant-coefficient derivatives, but it does not commute perfectly with the variable multiplication by $a(x)$. We obtain the precise equation for $v_{N}$:
$L(v_{N})=-P_{>N}(f_{k}(u))+\mathcal{C}_{N}(u_{t}),$ (35)
where $\mathcal{C}_{N}(u_{t})$ is the commutator between the spectral projector and the damping coefficient:
$\mathcal{C}_{N}(u_{t})=\operatorname{div}\Big([P_{>N},a(x)]\nabla u_{t}\Big).$ (36)
By the standard theory of pseudo-differential operators (see Appendix B for a detailed proof), the commutator of a zero-order multiplier ($P_{>N}$) with a smooth function $a\in W^{1,\infty}(\Omega)$ is a smoothing operator of order 1-1. Consequently, $P_{>N},a(x)]$ maps $L^{2}(\Omega)$ into $H^{1}(\Omega)$. The outer divergence operator then consumes exactly one spatial derivative, bringing the regularity gracefully back to $L^{2}(\Omega)$. Therefore, the commutator acts as a legitimate $L^{2}$-source bounded solely by the first-order energy, strictly independent of $k$:
$\|\mathcal{C}_{N}(u_{t})\|_{L^{2}(\Omega)}\leq C\|\nabla a\|_{L^{\infty}}\|\nabla u_{t}\|_{L^{2}(\Omega)}\leq C_{a}E_{0}^{1/2}.$ (37)
With a perfectly bounded right-hand side, we decompose the high-frequency component into its linear damped evolution and a perturbation: $v_{N}=v_{N}^{lin}+v_{N}^{per}$.
The linear part $v_{N}^{lin}$ is the solution to the homogeneous damped wave equation governed strictly by the high-frequency projection of the initial data:
$\begin{cases}L(v_{N}^{lin})=0\quad\text{(Homogeneous damped wave)}\\ v_{N}^{lin}(0)=P_{>N}u_{0}\\ \partial_{t}v_{N}^{lin}(0)=P_{>N}u_{1}.\end{cases}$ (38)
Since the localized dissipation only enhances the natural dispersive properties of the wave, $v_{N}^{lin}$ obeys the standard Strichartz estimates. Applying the estimate, we obtain:
$\|v_{N}^{lin}\|_{L^{5}(0,T;L^{10}(\Omega))}\leq C_{S}\left(\|\nabla(P_{>N}u_{0})\|_{L^{2}(\Omega)}+\|P_{>N}u_{1}\|_{L^{2}(\Omega)}\right)=:\epsilon(N).$ (39)
Since the initial data $(u_{0},u_{1})$ belongs to the finite energy space $H_{0}^{1}(\Omega)\times L^{2}(\Omega)$ and is completely independent of the truncation $k$, the sequence of partial sums of its spectral decomposition converges. Consequently, the high-frequency tail of the energy must vanish:
$\lim_{N\to\infty}\epsilon(N)=0.$ (40)
For the perturbed part $v_{N}^{per}$, which solves $L(v_{N}^{per})=-P_{>N}(f_{k}(u))+\mathcal{C}_{N}(u_{t})$ with zero initial data, we apply the inhomogeneous Strichartz estimate for the damped operator $L$:
$\|v_{N}^{per}\|_{L^{5}(0,T;L^{10}(\Omega))}\lesssim\|P_{>N}(f_{k}(u))\|_{L^{1}(0,T;L^{2}(\Omega))}+\|\mathcal{C}_{N}(u_{t})\|_{L^{1}(0,T;L^{2}(\Omega))}.$ (41)
Because the commutator source is uniformly bounded in space, its time integration yields:
$\|\mathcal{C}_{N}(u_{t})\|_{L^{1}(0,T;L^{2}(\Omega))}$ $\leq C\|\nabla a\|_{L^{\infty}}\|\nabla u_{t}\|_{L^{1}(0,T;L^{2}(\Omega))}$
$\leq C\cdot T\|\nabla a\|_{L^{\infty}}\|\nabla u_{t}\|_{L^{\infty}(0,T;L^{2}(\Omega))}$
$\leq C_{a}\cdot T\cdot E_{0}^{1/2}.$ (42)
Its contribution can therefore be made arbitrarily small by shrinking the time horizon $T$.
On the other hand, the nonlinear contribution is controlled inside the bootstrap regime:
$\|P_{>N}(f_{k}(u))\|_{L^{1}(0,T;L^{2}(\Omega))}\leq\|f_{k}(u)\|_{L^{1}(0,T;L^{2}(\Omega))}\lesssim\|u\|_{L^{5}(0,T;L^{10}(\Omega))}^{5}.$ (43)
Thus, the requisite smallness needed to close the high-frequency argument is obtained by combining: (i) the $T$-smallness of the commutator term, and (ii) the bootstrap control of the critical norm of $u$ on $(0,T)$. This yields:
$\|v_{N}\|_{L^{5}(0,T;L^{10}(\Omega))}\leq\epsilon(N)+C_{a}\cdot T\cdot E_{0}^{1/2}+C_{1}\|u\|_{L^{5}(0,T;L^{10}(\Omega))}^{5}.$ (44)
This mechanism provides the essential smallness required to absorb the critical quintic term, completely circumventing the need to assume that the total initial energy $E_{0}$ is small.